# The Value of AI Advice: Personalized and Value-Maximizing AI Advisors Are Necessary to Reliably Benefit Experts and Organizations

### v1

@@ Abstract
Despite advances in AI’s performance and interpretability, AI advisors can undermine experts’ decisions and increase the time and effort experts must invest to make decisions. Consequently, AI systems deployed in high-stakes settings often fail to consistently add value across contexts and can even diminish the value that experts alone provide. Beyond harm in specific domains, such outcomes impede progress in research and practice, underscoring the need to understand when and why different AI advisors add or diminish value. To bridge this gap, we stress the importance of assessing the value AI advice brings to real-world contexts when designing and evaluating AI advisors. Building on this perspective, we characterize key pillars – pathways through which AI advice impacts value – and develop a framework that incorporates these pillars to create reliable, personalized, and value-adding advisors. Our results highlight the need for system-level, value-driven development of AI advisors that advise selectively, are tailored to experts’ unique behaviors, and are optimized for context-specific trade-offs between decision improvements and advising costs. They also reveal how the lack of inclusion of these pillars in the design of AI advising systems may be contributing to the failures observed in practical applications.
@@ 1 Introduction
Artificial intelligence (AI) systems are poised to play an increasingly vital role in augmenting human decision-makers across high-stakes environments, such as in healthcare [1]. Yet, despite advances in AI’s generalization performance and interpretability, AI systems deployed to advise decision-makers in critical settings fail to consistently improve outcomes across contexts and expert behaviors [2, 3, 4, 5, 6, 7, 8]. For example, AI models with generalization accuracy that is comparable to or even exceeds that of human experts are increasingly used to advise radiologists in diagnostic tasks [5, 8, 3]. Yet, these tools have been shown to fall short of expectations, sometimes leading experts to make poorer decisions or prolonging the decision-making process [5, 9, 3] without adding sufficient value – outcomes which have also led to the under-utilization or abandonment of the tools after deployment [5].
The lack of benefit provided by AI tools is widespread across domains, task types, and human behaviors. A recent survey found that 56 papers on AI-assisted decision-making published between January 2020 and June 2023 reported that AI-assisted humans produced worse outcomes relative to the human alone [10]. Such detrimental outcomes run the risk of reducing trust in AI advice and impeding progress towards AI’s positive impact and adoption in critical decision-making domains. Nevertheless, to date, research on the development [10, 11, 12, 13] and evaluation [3, 14] of AI in AI-advised decision-making settings is often focused on improving AI’s state-of-the-art performance according to measures such as the AI’s accuracy on decision tasks without appropriate regard for how the AI’s advice will heterogeneously add to or diminish the value of the individuals it is meant to advise. To our knowledge, no existing work offers a comprehensive framework necessary to understand and mitigate the failures of AI advising systems seen in practice. Such a framework is critical to anticipating and explaining when AI advice will enhance or diminish the value provided by the individuals it aims to support. It is also necessary to guide the design of reliable, value-adding AI advisors. This paper seeks to address this gap.
We focus on a common high-stakes setting in which a human expert first makes an initial judgment on the best decision to take, and an AI may subsequently offer advice to the expert on the best decision to take before the expert finalizes their decision [5]. Henceforth we refer to this setting as AI-assisted Decision-Making (AIaDM), and refer to the AI as the AI advisor.
We propose that enhancing the positive impact of AI advisors in practice requires a fundamental shift toward a system-level, value-based approach to guide their design, learning, and evaluation. Consequently, the focus of this work is not the design of a new AI method with superior decision-making performance. First, we propose that AI advisors’ goal is to reliably offer the most added-value to their context, and we outline key pathways — referred to as pillars — through which AI advice can impact value in AIaDM contexts and are thereby integral to achieving the AI advisor’s goal. Building on the objective and pillars, we develop a framework that can be used to produce Reliable and Value-maximizing AI (ReV-AI) advisors, and we use our framework to develop a concrete proof-of-concept ReV-AI method – referred to as TeamRules (tr) – for generating inherently interpretable rule-based advisors, each personalized to an individual decision-maker and context. Using real world data along with simulation grounded in human behavior, we asses the value that different AI advisors’ provide across contexts, explore how different pillars affect AI advisors’ value impact, and shed light on the distinct advising strategies that ReV-AI advisors utilize to reliably add value.
We find that ReV-AI advisors produce consistent benefits across contexts and human behaviors by producing distinctly advantageous advising behaviors tailored to their context and advisee. Crucially, we also find that accounting for all of the pillars is key to yielding consistent benefits. Finally, our results reveal the risks posed by AI advisors which do not account for some of the pillars or elements of the framework, not only demonstrating how AI advisors can diminish value in some contexts, as has been shown to occur in practice, but also offering actionable insight into why such outcomes arise.
The AI Advising Objectives: Offer Reliable, Value-Maximizing Advice
@@
Perhaps the most fundamental objectives of an AI advisor are that (a) it learns to produce advice that maximizes the value it adds to its context relative to what can be achieved by the human alone and that (b) an AI advisor is reliable, such that when there is no opportunity to provide value-adding advice, it does not diminish the value delivered by the expert alone. Crucially for our work, the value added by an AI advisor encompasses not only the improvement in the decisions made by the human-AI team but also the practical costs associated with advising the human. Henceforth, we refer to advisors aiming to achieve these objectives as Reliable and Value-Maximizing AI (Rev-AI) advisors.
Pillars for Developing AI Advisors
@@
Central to our proposed system-level, value-based approach are pillars that characterize key pathways through which AI advice can add or diminish value in AIaDM settings and are thereby foundational for producing Rev-AI advisors and to our understandings of failures across applications in practice.
@@
  • The overlooked costs of AI advice. Recent studies on AI-advised classification tasks, such as medical diagnosis, often focus on optimizing the AI’s standalone predictive performance [11, 12, 13]. However, for AI advisors to add value in practice, improving predictive performance alone is insufficient. For instance, while double-reading diagnostic images by radiologists improves accuracy, it has not been widely adopted in the U.S. due to cost-effectiveness concerns [15, 16, 17]. Thus, to develop AI advisors that add value to their context, beyond improving decision outcomes, it is essential to account for the costs associated with integrating AI advice into the decision-making process.
@@
    A significant yet overlooked cost in the development and evaluation of AI advisors is the burden placed on experts when they engage with AI advice. AI advice can only improve an expert’s independent judgment when it contradicts it. Further, evidence shows that decision-makers in high-stakes settings experience significant time costs when reconciling AI advice that contradicts their judgment [18, 19, 9, 5]. Similar to previous innovations deemed cost-ineffective, excessive human engagement costs from reconciling disagreements with AI advice have been shown to impede the use of AI advice altogether [5, 20]. Advancing AI advisors’ positive impact on practice thus hinges on developing frameworks that produce cost-effective AI advisors, accounting for the human costs of reconciling contradictory advice. While crucial to the value proposition of AI advisors in critical settings, the human engagement and its implications have not been considered thus far in practice or in research on AI advising methods. Importantly, the value added by AI advice depends on these costs, and the cost-effectiveness of advice is moderated by the idiosyncratic trade-off within each context between added advising costs and the value generated from changes in the final decisions—which we henceforth refer to as the context’s trade-off.
  • Humans are idiosyncratic and imperfect advice-takers.
@@
    Human discretion in accepting or rejecting contradictory AI advice is both idiosyncratic and imperfect: individuals may reject accurate advice or accept incorrect recommendations [21, 22, 23, 24, 25, 26, 7, 8, 14]. We refer to this tendency as an individual’s Algorithm Discretion Behavior (ADB). ADB affects the cost-effectiveness of AI advice: given the human’s engagement to reconcile contradictory AI advice is costly, a higher likelihood of rejecting AI advice reduces its cost-effectiveness.
@@
  • Whether advice is accepted depends on the advice itself: offer inherently convincing advice.
@@
    An individual’s likelihood of accepting AI advice varies across decision instances and depends on properties inherent to the advice, such as its confidence [27, 28, 29]. Therefore, complementing AI recommendations with inherent properties of the advice has become crucial in high-risk settings, as such properties truthfully characterize the AI’s recommendation [30, 31]. We refer to inherent properties of AI advice that impact the likelihood with which an individual will accept it as ADB-bound information. Crucial for learning cost-effective advice, research shows that the human’s likelihood of accepting AI advice is influenced by the AI’s reported confidence in addition to the human’s self-reported (and potentially miscalibrated) confidence in their own decision [27, 28, 29]. Hence, developing value-maximizing AI advisors hinges on learning and offering advice that is simultaneously sufficiently inherently persuasive and accurate to add value, per the context’s trade-off.
@@
  • Humans are idiosyncratic and imperfect decision-makers. Lastly, the expert’s decision-making behavior is central to the development of cost-effective AI advisors. Due to the idiosyncratic strengths and weaknesses of experts’ judgment, the value of AI advice also depends on whether the AI’s recommendation is more likely to be correct than the human’s independent decision [32].
Framework for Producing Reliable Value-Maximizing AI Advisors
Figure 1: a) Framework for producing reliable, value-maximizing AI advisors. b) TeamRules + Human Partner Decision Making Process
@@
We build on the pillars to develop a general framework for producing ReV-AI advisors. In a departure from existing AI advisors, the ReV-AI framework constitutes a system-level and value-based approach for developing and evaluating AI advisors. ReV-AI advisors selectively offer advice that maximizes added value while avoiding actions that could diminish the value they provide to the expert within their context. Crucially, the advisors are grounded in the key pillars through which AI advice can impact the context’s value. Below we briefly note the key elements of the ReV-AI Learning Framework visualized in Figure 1a.
@@
Nature of advice: decision recommendation and inherent ADB-bound information. A ReV-AI advisor provides both a decision recommendation and ADB-bound information that influences the likelihood of its advice being accepted by the human advisee. We consider advisors that provide confidence—the estimated likelihood that the advice is correct—as the complementary ADB-bound information offered. However, our framework can also generate recommendations alongside other inherent properties that inform the human’s ADB and, by extension, the value of the advice. [^1] We focus on confidence here because it allows us to simulate diverse behaviors by building on prior work which has established with real human interactions how AI’s confidence along with the human’s self-assessed confidence affect the human’s receptiveness of advice [27].
@@
Learning to offer value-maximizing advice personalized to a given human and context. Our framework includes end-to-end learning of a model, which constitutes a repertoire of decision recommendations with corresponding confidence levels that are offered selectively during deployment, so as to add the most value across the decision space. Note that to add the most value to a context, a ReV-AI advisor can learn to produce advice that adds significant value for some decisions at the possible expense of accuracy or persuasiveness on other decisions. Value-maximizing advice is learned through an objective grounded in the pillars, which reflect how AI counsel either enhances or diminishes value in its context. The value of the advice is thus influenced by (1) its inherent persuasiveness given the human’s confidence and corresponding ADB, (2) its likelihood of being correct relative to the likelihood the human’s independent decision is correct, and (3) the context’s trade-off between the human’s engagement with the AI recommendation and the decision benefits of the advice. Hence, producing value-maximizing advice grounded in the ReV-AI advising objective requires data on the human’s expertise, confidence, and ADB. We discuss data collection and related modeling in the Methods section.

### v2

@@ Abstract
Despite advances in AI’s performance and interpretability, AI advisors can undermine experts’ decisions and increase the time and effort experts must invest to make decisions. Consequently, AI systems deployed in high-stakes settings often fail to consistently add value across experts and organizations and can even diminish the value that experts alone provide. Beyond harm in specific domains, such outcomes impede progress in research and practice, underscoring the need to understand when and why different AI advisors add or diminish value. To bridge this gap, we stress the importance of assessing the value AI advice brings to real-world contexts when designing and evaluating AI advisors. Building on this perspective, we characterize key pillars – pathways through which AI advice impacts value – and develop a framework that incorporates these pillars to create reliable, personalized, and value-adding advisors. Our results highlight the need for value-driven development of AI advisors that advise selectively, are tailored to experts’ unique behaviors, and are optimized for context-specific trade-offs between decision improvements and advising costs. They also reveal how the lack of inclusion of these pillars in the design of AI advising systems may be contributing to the failures observed in practical applications.
@@ 1 Introduction
Artificial intelligence (AI) systems are poised to play an increasingly vital role in augmenting human decision-makers across high-stakes environments, such as in healthcare (pnas_occupation_2024, ). Yet, despite advances in AI’s generalization performance and interpretability, AI systems deployed to advise decision-makers in critical settings fail to consistently improve outcomes across contexts and expert behaviors (jama_2024, ; Heterogeneity_Salz_Agarwal_Rajpurkar_2024, ; green_disparate_2019, ; lebovitz_engage_2022, ; green_algorithmic_2021, ; Jacobs2021-zs, ; Rezazade_Mehrizi2023-ia, ). For example, AI models used to advise radiologists on diagnostic tasks (lebovitz_engage_2022, ; Rezazade_Mehrizi2023-ia, ; Heterogeneity_Salz_Agarwal_Rajpurkar_2024, ) and having generalization accuracy comparable to or exceeding that of human experts have failed to consistently add value, sometimes leading experts to make poorer decisions or prolonging the decision-making process without adding sufficient value (lebovitz_engage_2022, ; shinradio2023, ; Heterogeneity_Salz_Agarwal_Rajpurkar_2024, ). These outcomes have also contributed to the under-utilization or abandonment of the tools after deployment (lebovitz_engage_2022, ). More broadly, a recent survey found that 56 papers on AI-assisted decision-making published between January 2020 and June 2023 reported that AI-assisted humans produced worse outcomes relative to what they were capable of alone (Vaccaro2024, ). Such detrimental outcomes not only run the risk of reducing trust in AI advice but impeding progress towards AI’s positive impact and adoption in critical decision-making domains.
Despite AI advising systems’ failures to consistently add value in practice, prior work on methods for producing AI advisors has not offered a comprehensive framework for producing AI advising systems that reliably add value and mitigate losses seen in practice. Such a framework is crucial to guide the design and evaluation of reliable, value-adding AI advisors as well as to facilitate AI’s positive impact in an array of critical decision-making domains. This paper seeks to address this gap.
Figure 1: AI-Advised Decision Making
We focus on high-stakes settings that follow an AI-advised decision-making paradigm common in practice, shown in Figure 1, where an expert first makes an independent judgment on the best decision to take, and an AI subsequently advises the expert on the decision before the expert finalizes their decision (lebovitz_engage_2022, ; GOMEZ2023102977, ; Goh2025, ; Yin2025, ; when_advise_2023, ; yin2019, ; riverain2022clearread, ). [^1] As an example, ClearRead Xray, used by hundreds of institutions in 13 countries, offers diagnostic advice to radiologists after the physician has performed an initial interpretation of the radiograph. For brevity, throughout this paper we refer to such settings as AI-advised Decision-Making (AIaDM) settings and we refer to the AI in AIaDM settings as the AI advisor. We focus on such settings because of their practical impact; however, the process and principles that we propose are general and can be adapted to produce value-based and reliable AI partners across contexts with different forms of human-AI collaboration.
We propose that enhancing the positive impact of AI advisors in practice requires a shift towards a value-based approach to guide AI advisors’ design, learning, and evaluation. Foundational to the framework we develop is a design that builds on opportunities to produce AI advisors that add the most value to their contexts and that simultaneously mitigate the losses they might introduce. Despite the failures in practice, prior research on the development (pnas_ecg, ; Nishida2023, ; Abbas2024, ) and evaluation (Heterogeneity_Salz_Agarwal_Rajpurkar_2024, ; NBERw31422, ; Vaccaro2024, ) of AI advisors has not proposed a framework for developing AI advisors that engage with and account for the different pathways through which AI advice may heterogeneously either add or diminish value across contexts. Furthermore, while AI advisors in practice are trained to produce the most accurate decisions if the AI were to make these decisions autonomously, we consider settings in which the AI does not undertake decisions autonomously and may only advise a human, thus our design engages with how AI advice may impact the human’s decision outcomes.
To our knowledge, this work is the first to (1) propose a value-based framework for producing AI advisors in AIaDM contexts, (2) empirically study its performance across diverse contexts relative to alternatives, (3) identify and enhance our understanding of contexts in which different AI advisors can add the most value or risk the greatest losses, and (4) uncover insights on the unique expertise and advising behaviors that AI advisors can give rise to and how they contribute to AI advisor’s added value in their contexts. More concretely, our contributions are as follows:
  1. 1.
    We first propose that the design objective of AI advisors is to reliably offer the most added-value to their context, which entails accounting for pathways through which AI advice can impact decision outcomes but also considering the costs arising from the human’s engagement with AI advice.
  2. 2.
    Foundational to our contributions is the identification of key pathways — referred to as pillars — through which AI advice can impact the value added or diminished by an AI advisor and are thereby integral to achieving the AI advisor’s goal. To the best of our knowledge, no prior work that developed AI advisors has proposed and motivated the benefits of simultaneously accounting for these pillars, and proposed methods for producing AI advisors that account for them so as to add the most value. Importantly, the pillars’ effects on decision-making outcomes are interdependent and thus introduce novel understandings of how AI advisors can add value or might lead to losses.
  3. 3.
    Building on the objective and pillars, we propose a framework for producing Reliable and Value-Maximizing AI (ReV-AI) advisors. ReV-AI advisors’ expertise and advising behaviors are shaped by their context: they aim to complement the individual advisee’s expertise and advice-taking behaviors, while accounting for the context’s tradeoff between the costs and potential benefits from AI advice. ReV-AI advisors selectively offer advice and are specialized to learn and offer advice that is simultaneously sufficiently accurate and inherently convincing to the advisee so as to add the most value.
  4. 4.
    We use our framework to develop, evaluate, and analyze a concrete proof-of-concept ReV-AI method – referred to as TeamRules (tr) – for generating inherently interpretable (ReV-AI) advisors.
  5. 5.
    Using real world data along with simulation grounded in human behavior, we explore the value that different AI advisors’ can add or diminish across diverse contexts corresponding to different data domains, differential human decision-making abilities, advice-taking behaviors, and organizational trade-offs between decision benefits and costs. Our results reveal that across contexts, human expertise and advice-taking behaviors, ReV-AI advisors add significantly greater value and otherwise avoid diminishing the value offered by the human alone. Overall, our results demonstrate that producing AI advisors that can be relied on in practice to add the most value and avoid losses in high-stakes contexts requires that they offer effective responses by accounting for the pathways through which AI advice can add or diminish value. Consistent with evidence from practice, we also find that the most prevalent AI advisors in practice, not only may not add the most value, but introduce the greatest losses relative to the value already offered by the human alone. We also show why these outcomes may arise and reveal the contexts in which the most prevalent AI advisors can yield the greatest losses.
  6. 6.
    We analyze two case studies grounded in real-world expert decision-making and advice-taking behavior to explore the expertise and advising strategies that ReV-AI advisors give rise to relative to alternatives. Our analysis reveals that ReV-AI advisors produce distinct expertise and advising behaviors that are shaped by their context and advisee and that these behaviors contribute directly to greater value creation and loss mitigation.
Our work advocates and aims to build the groundwork for value-based AI advisors in high-stakes settings. We hope that our work will inspire future research to advance AI’s positive impact on high-stakes decisions and to develop other value-based AI frameworks for different forms of human-AI collaborations and contexts. Our work also aims to equip managers and stakeholders with a clearer understanding of the opportunity costs and limitations of current AI advisory systems and the contexts in which these limitations are likely to be most pronounced, while offering guidance to advance next-generation AI advisors that reliably enhance decision-making and consistently drive positive real-world outcomes.
2 The Case for Reliable & Value-Driven AI Advisors
In this section, we discuss the goal of AI advisors and build on prior work to introduce pathways through which advice impacts value. We then introduce a framework for developing AI advisors that reliably add value by accounting for these pathways.
2.1 The AI Advising Objectives: Offer Reliable, Value-Maximizing Advice
@@
Perhaps the most fundamental objectives of an AI advisor are that (a) it learns to produce advice that maximizes the value it adds to its context relative to what can be achieved by the human alone and that (b) it is reliable, such that when there is no opportunity to provide value-adding advice, it does not diminish the value delivered by the expert alone. Henceforth, we refer to advisors aiming to achieve these objectives as Reliable and Value-Maximizing AI (ReV-AI) advisors. Crucially for our work, the value added by an AI advisor encompasses not only the improvement in the decisions made by the human-AI team but also the practical costs associated with advising the human. Thus, developing AI advisors that effectively add value does not correspond to developing models that have high standalone performance on a given task; rather, AI advising models need to cost-effectively improve an expert’s decision outcomes and avoid diminishing them. Notably, this means that a worse-performing model that complements an individual’s weaknesses while leveraging their strengths may be a better AI advisor to that expert than a model that has higher standalone performance across the decision space.
2.2 Pillars for Developing AI Advisors
@@
Central to our proposed value-based approach are pillars that characterize key pathways through which AI advice can add or diminish value in AIaDM settings and are thereby foundational to producing reliable and value-maximizing AI advisors.
2.2.1 The overlooked costs of AI advice.
@@
Recent studies on AI-advised classification tasks, such as medical diagnosis, often focus on optimizing the AI’s standalone predictive performance (pnas_ecg, ; Nishida2023, ; Abbas2024, ). However, for AI advisors to add value in practice, improving predictive performance alone is not always sufficient. For instance, while double-reading diagnostic images by radiologists improves accuracy, it has not been widely adopted in the U.S. due to cost-effectiveness concerns (stefano_double_1995, ; Posso2016, ; pmid29594850, ). Concretely, given double-readings double the diagnostic costs, its integration in practice hinges on whether the benefits in outcomes outweigh the costs. Consequently, in many key high-stakes contexts, for AI advisors to add value it is essential to account for any added costs incurred when AI advice is integrated into the decision-making process.
@@
A significant cost, which has been largely overlooked in the development and evaluation of AI advisors, is the burden placed on experts when they engage with AI advice that conflicts with their independent judgment. AI advice can only improve an expert’s independent judgment when it contradicts it. Yet, evidence shows that decision-makers in high-stakes settings experience significant time costs when reconciling AI advice that contradicts their judgment (Reverberi2022, ; Boyac2023, ; shinradio2023, ; lebovitz_engage_2022, ). Similarly to previous innovations deemed cost-ineffective, excessive human engagement costs from reconciling disagreements with AI advice have been shown to impede the use of AI advice altogether (lebovitz_engage_2022, ; Lenharo2024, ).
Advancing AI advisors’ positive impact on practice thus hinges on developing frameworks that produce cost-effective AI advisors by accounting for the human costs of reconciling contradictory advice. While crucial to the value proposition of AI advisors in critical settings, the human engagement and its implications have not been considered thus far in practice or in research on AI advising methods. Importantly, the value added by AI advice depends on these costs, and the cost-effectiveness of advice is moderated by the trade-off within each context between added advising costs and the value produced from changes in the final decisions—which we henceforth refer to as the context’s trade-off. The framework we propose below accounts for the advising costs incurred to yield AI advisors that consistently add value. However, as we also demonstrate in the empirical evaluations, the framework generalizes to contexts which allow any added costs as long as the advice improves decision outcomes.
2.2.2 Humans are idiosyncratic and imperfect advice-takers.
@@
Human discretion in accepting or rejecting contradictory AI advice is both idiosyncratic and imperfect: individuals may reject accurate advice or accept incorrect recommendations (dietvorst_algorithm_2015, ; green_disparate, ; green_principles, ; stop, ; bansal_does_2021, ; complete-me, ; Jacobs2021-zs, ; Rezazade_Mehrizi2023-ia, ; NBERw31422, ). Prior work has shown that decision-makers tendency to accept and reject advice is informed by the human’s own (and, possibly, miscalibrated) confidence as well as the AI’s confidence in a given decision, both of which may vary across decision instances and differ across different decision-makers (will_you_accept, ; CHONG2022107018, ). We refer to this tendency as an individual’s Algorithm Discretion Behavior (ADB). Prior work has also shown that humans’ ADB is imperfect and they consequently do not necessarily benefit from AI advice even if the AI offers reliability measures such as confidence or is explainable (Vaccaro2024, ; bansal_does_2021, ). Importantly, ADB affects the cost-effectiveness of AI advice: given the human’s engagement to reconcile contradictory AI advice is costly, a higher likelihood of rejecting AI advice reduces its cost-effectiveness.
@@
2.2.3 Whether advice is accepted depends on the advice itself: offer inherently convincing advice.
@@
An individual’s likelihood of accepting AI advice varies across decision instances and depends on properties inherent to the advice, such as its confidence (will_you_accept, ; CHONG2022107018, ; zhang_effect_of_conf, ). Therefore, complementing AI recommendations with inherent properties of the advice has become crucial in high-risk settings, as such properties truthfully characterize the AI’s recommendation (laugel2019, ; rudin2019stop, ). We refer to inherent properties of AI advice that impact the likelihood with which an individual will accept it as ADB-bound information. Crucial for learning cost-effective advice, research shows that the human’s likelihood of accepting AI advice is influenced by the AI’s reported confidence [^2] While humans’ likelihood of accepting AI advice is influenced by AI’s reported confidence, this influence does not necessarily lead to appropriate likelihood of accepting AI advice. Thus, as mentioned in Section 2.2.2, humans’ may accept bad advice and reject good advice even if the AI offers reliability measures such as confidence. in addition to the human’s self-reported (and potentially miscalibrated) confidence in their own decision (will_you_accept, ; CHONG2022107018, ; zhang_effect_of_conf, ). Hence, developing value-maximizing AI advisors hinges on learning and offering advice that is simultaneously sufficiently inherently persuasive and accurate to add value, per the context’s trade-off.
2.2.4 Humans are idiosyncratic and imperfect decision-makers.
@@
Lastly, the expert’s decision-making behavior is central to the development of cost-effective AI advisors. Due to the idiosyncratic strengths and weaknesses of experts’ judgment, the value of AI advice also depends on whether the AI’s recommendation is more likely to be correct than the human’s independent decision (bansal_is_2021, ). Importantly, given AI is an artifact that can be intentionally designed to add the most value, different experts’ idiosyncratic strengths and weaknesses introduce opportunities to produce AI advisors with differential strengths to complement them.
2.3 Framework for Producing Reliable Value-Maximizing AI Advisors
@@
We build on the pillars to develop a general framework for producing ReV-AI advisors. In a departure from existing AI advisors, the ReV-AI framework constitutes a value-based approach for developing and evaluating AI advisors. ReV-AI advisors selectively offer advice that maximizes added value while avoiding actions that could diminish the value they provide to the expert within their context. Crucially, the advisors are grounded in the key pillars through which AI advice can impact the context’s value. Below we briefly note the key elements of the ReV-AI Learning Framework visualized in Figure 2 and we formalize the framework in Section 2.4.
Figure 2: Framework for producing reliable, value-maximizing AI advisors
@@
Nature of advice: decision recommendation and inherent ADB-bound information. A ReV-AI advisor provides both a decision recommendation and ADB-bound information that influences the likelihood of its advice being accepted by the human advisee. We consider advisors that provide confidence—the estimated likelihood that the advice is correct—as the complementary ADB-bound information offered. However, our framework can also be used to generate recommendations complemented by other inherent properties of AI advice that impact the human’s ADB and, by extension, the value of the advice. We focus on confidence here because prior work has established how AI’s reported confidence along with the human’s self-reported confidence affect the human’s receptiveness of advice (will_you_accept, ), allowing us to explore outcomes in diverse simulated AIaDM contexts.
@@
Learning to offer value-maximizing advice personalized to a given human and context. Our framework includes end-to-end learning of a model, which constitutes a repertoire of decision recommendations with corresponding confidence levels that are offered selectively during deployment, so as to add the most value across the decision space. Note that to add the most value to a context, a ReV-AI advisor can learn to produce advice that complements the human’s expertise by adding significant value for some decisions that the human can benefit the most from at the possible expense of accuracy or persuasiveness on other decisions. Value-maximizing advice is learned through an objective grounded in the pillars, which reflect how AI counsel either enhances or diminishes value in its context. The value of the advice is thus influenced by (1) its inherent persuasiveness given the human’s confidence and corresponding ADB, (2) its likelihood of being correct relative to the likelihood the human’s independent decision is correct, and (3) the context’s trade-off between the human’s engagement with the AI recommendation and the decision benefits of the advice. Hence, producing value-maximizing advice grounded in the ReV-AI advising objective requires data on the human’s expertise, confidence, and ADB. We discuss data collection and related modeling in Section 3.